OpenStudy (anonymous):

What is the quotient of c^2-c-2/c^2-1 / c^2+3c+2/c^2-3c+2 in lowest terms? I am completely lost on this... :(

OpenStudy (johnweldon1993):

\[\huge \frac{c^2 - c - 2}{\frac{c^2 - 1}{\frac{c^2 + 3c + 2}{c^2 - 3c + 2}}}\]

OpenStudy (anonymous):

more or less yes that's what it's asking. I need to bring the quotient of the equation to the lowest terms... (That is if I'm understanding what you're saying)

OpenStudy (johnweldon1993):

Well when you are dividing by a fraction...you take the bottom fraction...flip it...and turn this into multiplication... so all that up there...becomes... \[\huge \frac{c^2 - c - 2}{c^2 - 1} \times \frac{c^2 - 3c + 2}{c^2 + 3c + 2}\]

OpenStudy (johnweldon1993):

See things that cancel?

OpenStudy (anonymous):